HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SYSTEMATIC INCREASE OF WATER VIBRATIONAL EXCITATION IN AR–WATER VAN DER WAALS COMPLEX

\underline{R. GLORIEUX} , Institute of Condensed Matter and Nanosciences, Université catholique de Louvain, Chemin du cyclotron, 2 1348 Louvain-la-Neuve, Belgium.;

N. MOAZZEN-AHMADI, A. J. BARCLAY, Department of Physics and Astronomy, University of Calgary, 2500 University Drive North West, Calagry, Alberta T2N 1N4, Canada.; M. HERMAN, T. VANFLETEREN, Spectroscopy, Quantum Chemistry and Atmospheric Remote Sensing (SQUARES), Faculté des Sciences, Université libre de Bruxelles (ULB), 50 ave. F-D Roosevelt, B-1050,

Brussels, Belgium.; A. R. W. MCKELLAR, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.; B. M. HAYS, Université de Lille, CNRS, UMR8523-PhLAM-Physique des Lasers Atomes et Molécules, F-59000 Lille, France.; A. S. BOGOMOLOV, C. LAUZIN, Institute of Condensed Matter and Nanosciences, Université catholique de Louvain, Chemin du cyclotron, 2 1348

Louvain-la-Neuve, Belgium.

The Ar–water van der Waals complex is one of the most studied complexes. This work focuses on new results obtained in the mid- and near-infrared. Complexes were formed in a pulsed slit supersonic jet expansion and were probed using sensitive laserbased absorption spectrometers. In this talk, I will present the rotationally resolved spectra of the Ar–water as a function of the number of vibrational quanta of excitation in the water molecule. Note that the dissociation energy of Ar–D_{2}O is evaluated to be below 100 cm^{−1} (D_{0} ∼ 95 cm^{−1} ∼ 12 meV).^{1} We started by probing the 1OD stretch region of Ar–D_{2}O, i.e. the (ν′_{1}, ν′_{2}, ν′_{3}) ← (ν′′_{1}, ν′′_{2}, ν′′_{3}) = (0, 0, 1) ← (0, 0, 0) vibrational band, where ν_{1}, ν_{2}, and ν_{3} are the vibrational quantum numbers of the isolated water molecule. This range is close to 30 times the dissociation energy. The spectra were assigned and observed perturbations are discussed following previous study on Ar–H_{2}O.^{2} After making improvements on the FANTASIO experimental setup,^{3} we observed new bands in the 2OH stretch region of Ar–H_{2}O.^{4}^{,}^{5} In this region, the excitation energy is approximately 77 times D_{0}.^{6} Finally, preliminary results concerning the 3OD stretch excitation in Ar–D_{2}O (83 times D_{0}) will be presented.

1S. Wang, S. He, L. Dai, E. Feng, and W. Huang, J. Chem. Phys. 142, 224307 (2015).

2D. J. Nesbitt, and R. Lascola, J. Chem. Phys. 97, 8096 (1992).

3A. S. Bogomolov, A. Roucou, R. Bejjani, M. Herman, N. Moazzen-Ahmadi, and C. Lauzin, Chem.

Phys. Lett. 774, 138606 (2021).

4T. Vanfleteren, T. Földes, M. Herman, Chem. Phys. Lett. 627, 36-38 (2015).

5T. Vanfleteren, T. Földes, M. Herman, J. Liévin, J. Loreau, and L. H. Coudert, J. Chem. Phys. 147, 014302 (2017).

6D. Hou, Y.-T. Ma, X.-L. Zhang, and H. Li, J. Chem. Phys. 144, 014301 (2016).

p-number: p138  Submitted on Wed, 14 Jun 2023 16:26:51 +0200